The isotope strontium-90 is produced during the testing of nuclear weapons. If 100.0mg of strontium-90 was released in the atmosphere in 1960, how much of the radioisotope remains 85 years later? The half-life of strontium-90 is 29 years. Show all your work.

1
Expert's answer
2022-06-07T09:45:03-0400

Solution:

Strontium-90 ⇨ Sr-90

The half-life (t1/2) of strontium-90 is 29 years


From law of radioactive decay:

t1/2 = ln(2) / λ

where λ = decay constant

λ = ln(2) / t1/2

λ = (0.693) / (29 years) = 0.0239 years−1


Radioactive decay follows the kinetics of a first order reaction:

ln[Sr-90]t = −λt + ln[Sr-90]o

 

λ = 0.0239 years−1

t = 85 years

[Sr-90]o = 100.0 mg

[Sr-90]t = unknown


Therefore,

ln[Sr-90]t = −(0.0239 years−1 × 85 years) + ln(100.0)

ln[Sr-90]t = −2.0315 + 4.60517 = 2.57367

[Sr-90]t = e2.57367 = 13.11

[Sr-90]t = 13.11 mg


Answer: 13.11 mg of the radioisotope (Sr-90) remained